# MOND as a transformation between non-inertial reference frames via Sciama’s interpretation of Mach’s Principle

### v3

@@ 3 Discussion
It is well known that, according to Bekenstein, a higher value for Milgrom’s MOND $a_{0}$ could resolve the tension between Milgrom’s MOND and galaxy cluster’s core dynamics (and that $a_{0}$ seems to grow with scale), where Newtonian accelerations take around that same value (provided that $a_{0}$ stays the same in galaxies for agreement with rotation curves). Machian MOND, although being equivalent to Milgrom’s MOND, is effectively a model of a varying $a_{0}$ with time scale and could in principle help resolving this issue of Milgrom’s MOND. It could also in principle solve the cosmological issues of Milgrom’s MOND [35]. As modified gravity, the equivalent variation of the Newtonian $G$ is both temporal due to the expanding universe and dependence on the mass and radius of the observable universe, and spatial since local potentials and local field intensities should also be included in Sciama’s relationship and the Machian MOND transformation, respectively.
@@
The form of the transformation (4) is not unique and depends on the chosen Milgrom’s MOND interpolating function for which the Machian substitutions are made. The exact form of the transformation should not be fundamental but derived as an approximation from a complete FundaMOND theory of modified inertia considering the relativity of acceleration and inertia. Furthermore, which integers and mathematical constants play a role within the Machian transformation and within Sciama’s relationship is left to be discussed for a better agreement with observations.
@@ Acknowledgments
The authors acknowledge a similar but not exactly the same as that of equation (4) to Juan D. Santander.

### v4

@@ 3 Discussion
It is well known that, according to Bekenstein, a higher value for $a_{0}$ could resolve the tension between Milgrom’s MOND and galaxy cluster’s core dynamics (and that $a_{0}$ seems to grow with scale), where Newtonian accelerations take around that same value (provided that $a_{0}$ stays the same in galaxies for agreement with rotation curves). Machian MOND, although being equivalent to Milgrom’s MOND, is effectively a model of a varying $a_{0}$ with time scale and could in principle help resolving this issue of Milgrom’s MOND. It could also in principle solve the cosmological issues of Milgrom’s MOND [38], such as those arising from a fixed scale-dependence on $a_{0}$. As modified gravity, the equivalent variation of the Newtonian $G$ is both temporal due to the expanding universe and dependence on the mass and radius of the observable universe, and spatial since local potentials and local field intensities should also be included in Sciama’s relationship and the Machian MOND transformation, respectively.
@@
The form of the transformation (4) is not unique and depends on the chosen Milgrom’s MOND interpolating function for which the Machian substitutions are made. Although (4) might be in conflict with solar system constraints, as it approaches $\mu(x)\rightarrow 1$ too slowly at the high acceleration regime, this is a problem of the slope of the transformation, which can be chosen to be different by another form for it. The exact form of the transformation should not be fundamental but derived as an approximation from a complete FundaMOND theory of modified inertia á la Mach considering the relativity of acceleration and inertia. Furthermore, which integers and mathematical constants play a role within the Machian transformation and within Sciama’s relationship is left to be discussed for a better agreement with observations.
@@ Acknowledgments
The authors acknowledge a similar but not exactly the same as that of equation (4) to Juan D. Santander. We also thank Mordehai Milgrom for his warm and helpful feedback.